OpenStudy (anonymous):

How do I find the maximum value of: y = -3x^2 - 9x + 2 How do I find the maximum value of: y = -3x^2 - 9x + 2 @Algebra

OpenStudy (anonymous):

using calculus ... find the derivative and set it = 0 not using calculus fund the vertex

OpenStudy (anonymous):

Its for an Intermediate Algebra course haha

OpenStudy (anonymous):

f'(x)= -6x -9

OpenStudy (anonymous):

ok determine the vertex y=-3(x^2+3x) +2 y=-3(x^2+3x +9/4-9/4) +2 y=-3(x +3/2)^2 +27/4+2 y=-3(x+3/2) +35/4 so max is 35/4 ... but check my arithmetic

OpenStudy (anonymous):

max vale is the y value for the vertex vertex is at (-3/2,35/4)

OpenStudy (anonymous):

So basically yuou just completed the square?

OpenStudy (anonymous):

yes ... it was completed ... poor old thing ...

OpenStudy (anonymous):

Thanks :)